2. The difference between two numbers is 28. The larger number is four more than thrice the smaller.
What are the numbers?

Answer by dileepkumarabhijith(12)

hello
let the numbers be x and y
diff between them (X-Y)=28.........(1)
clearly larger number is x
X=4+3y..............(2)
substituting equation 2 in (1)-------> 4+3y-y=28
4+2y=28 2y=24
y=12
x can be found out substituting value of y in(1)
that is x-12=28
therefore, x=28+12=40